Distance traveled by Hamid and Akhtar in an hour is 9 km and 12 km. Find the ratio of the speed of Hamid to the speed of Akhtar

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We know that the speed of a certain object is the distance travelled by that object in an hour Distance travelled by Hamid in one hour = 9 km
Distance travelled by Akhtar in one hour = 12 km
Speed of Hamid = 9 km/hr
Speed of Akhtar = 12 km/hr
Ratio of speed of Hamid to the speed of Akhtar = 9 / 12 = 3 / 4

Thanks